From GlueXWiki
Jump to: navigation, search

June 7, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
    • Production status
    • Fourth package tests (Vlad)
  2. Engineering update (Bill)
  3. Electronics update (Chris)
  4. Gas system (Beni)
  5. Test set-up in 126
  6. Other FDC E-log


Participants: Bill, Dave, Chris, Nick, Vlad, Beni, and Lubomir.


- Dave: working on the last wire frame, Casey finished wire gluing. Chris will put elements on a wire frame tomorrow. Also two wire planes will be deadened tomorrow/Monday. Cathode #47 (type-2) had a problem on strip #1: there is a small bump of epoxy on the trace leading to the connector; it was put there before to fix small deformation of the Kapton foil. Since the bump was very close to the O-ring and solid g10 surfaces we were afraid that it may create gas leakage. After trying to remove part of the epoxy we damaged the copper trace. At this point we decided to put another cathode in the fourth package. Later the trace was fixed by applying conductive epoxy (3-4 Ohms strip to connector resistance) but we decided to use this cathode as a spare and make another cathode for the fourth package (#55) which will be ready by Wednesday-Thursday. Also by that time we will have all the wire frames ready to finish the fourth package.

- Dave predicts we will have all the elements, also the spares, done by the end of June. Then the work will be focused on the first and second package refurbishment. Related to that Bill asked when we can do full package tests in 126. Answer: everything depends on the fADC125 availability. We need to discuss this with Fernando since it will delay the FDC production/testing, and make a plan at least for partial testing of the cathodes using as many fADC125 as possible.

- Three cells were installed on the fourth package at the end of the last week. Vlad reported on the tests: oxygen went down to 77ppm, all channels were tested with scope/source and are working fine. However, on the top cell (#4) he found a point source of noise (~50kHz, high amplitude) visible on one wire (#21), top strip and bottom strip; most likely standard hair-on-wire problem. On cell #5 Vlad found that the current increased in one day from ~20nA to ~200nA and then it was stable. We will keep the HV on #5 and depending how stable it is we will decide if we want to open it as well, or open ONLY the top cell (#4) to fox the high frequency noise.

Engineering update

- Bill: busy with the tagger, but works with Caleb on the cooling system. Related to a question about the inter-package support tubes, Bill explained that the cables will not be tightened to that tube. This is because the diameter is smaller, the cables would be further away from BCAL and reduce the photon efficiency. To keep the cables straight between the packages Bill propose to use plastic rods/tubes inserted in the cable bundles.

Electronics update

- Chris: grounding clips will come end of next week. Then we start with the ground installation on the third package at Blue Crab; Beni will help us. We discussed what additional material we will use: 1mil copper foil for grounding strips, Rohacell wrapped with something (Al foil?), and clips that we already have.

- Chris got a quote from San Francisco Circuits about purchase of additional PCBs, in case we can't fix some of the first four wire frames: $26.9K for 4 sets, and $37K for 8 sets (each set has 6 PCBs). These are quotes only the board production; Chris can do the stuffing of the boards. We will replace the caps on the first wire frame and then decide if we need extra PCBs.

Gas system

- Beni is in contact with Narciso Gomez about the gas system. We discussed if we would need to filter the water that according to the gas company would be less than 200ppm in the CO2 bottles. Lubomir: don't have to filter since 200ppm is low enough especially when the oxygen level is low. Beni was advised to contact the GTS gas company since JLab has a contract with it.

Test set-up at 126

- Vlad will start testing all the cathode channels of the first package with a source and scope. We want to see if there was some deterioration of the cathode connections done with the z-axis conductive tape. There are some indications for increased resistance (up to 5-6 Ohms) on a small percentage of cathode channels, about 3-4 channels per cathode, at Blue Crab. This, however, could be due to handling of the cathodes and once they are installed in a package there's no deterioration.


- Lubomir showed a GEANT model of the Helium Drift Chamber (at the FDC E-log). The model will be used first to study the BCAL photon efficiency. Important is that the diameter of the solid frame was increased and is now closer to the BCAL. This is possible because one can put the readout cards parallel to the z-axiz; this was not possible for the FDC due to the high density of the channels.